dimension of magnetic induction

Grade:12th Pass

5 Answers

Digant Dubey
12 years ago

M T -2 A -1

Nikhil Cherukupalli
12 years ago

We know that F= B I L... From that we get B as F/IL = [MLT-2] / [A] [L] = [ML0T-2A-1]

Anjali kumari
5 years ago
It is equals to Magnetic field intensity vector which is equals to MT^-2A^-1.
Prabhavathi Siram
5 years ago
From the equation F= BIL ,
B= F/IL
B=magnetic induction
I= Current (in Amperes)
F= Force=ma
L= length=l
B= ma/(amp.*l)
B= MLT^-2/(A*L)
   =ML°T^-2 A^-1
 
 
L=length
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

From the equation F= BIL ,
B= F/IL
B=magnetic induction
I= Current (in Amperes)
F= Force=ma
L= length=l
B= ma/(amp.*l)
B= MLT^-2/(A*L)
=ML°T^-2 A^-1

Thanks and Regards